# Byzantine-Eavesdropper Alliance: How to Achieve Symmetric Privacy in Quantum $X$-Secure $B$-Byzantine $E$-Eavesdropped $U$-Unresponsive $T$-Colluding PIR?

### v1

@@ 1.2 Related Work
As for the QPIR literature, in [7, 35] the eavesdropper variation was studied and the rate regions were found. In [36], the quantum unresponsive servers variation was analyzed and the rate regions were given. Finally, the quantum Byzantine servers version was analyzed and the rate regions were found in [37]. More applications for the new $N$-sum box abstraction can be found in [38, 39, 40, 41].
@@ 2 Problem Formulation
A user chooses $\theta$ uniformly at random from $[K]$ to retrieve the $\theta$th message. The user sends the query $Q_{n}^{[\theta]}$ over a classical channel to the $n$th server. Any $\mathcal{T}\subset[N]$ servers can collude to attempt to decode the required message index $\theta$, where $|\mathcal{T}|\leq T$. Thus, the queries must be secure against any $\mathcal{T}$ colluding servers, i.e.,
@@
Upon receiving all the answers, the user must be able to decode the required message $W_{\theta}$, i.e.,
$S(W_{\theta}|\mathcal{A}_{[N]},Q_{[N]})=0,$ (9)
where $S(\cdot)$ is the Von Neumann quantum entropy.
@@ 3 Main Results
[…]
@@
[…]
@@
In Theorem 2, the security parameter changes from $X$ to $H=\max\{X,B\}$. Intuitively, the main reason behind this is to make sure that, if all $B$ Byzantine servers send the $m$th symbol for the $\kappa$th message $W_{\kappa,m}$ to the user, the user will not be able to decode it.
@@ 4.1 Quantum Information Theory
For a quantum system $A$, that can be in the state |$\ket{\psi_{j}}$ with probability $p_{j}$, the quantum density matrix $\rho_{A}$ is defined as,
$\mathtt{\rho}_{A}=\sum_{j}p_{j}\ket{\psi_{j}}\bra{\psi_{j}},$ (15)
@@ 4.2 $N$-Sum Box Abstraction
The encoding and decoding structure using the $N$-sum box abstraction introduced recently in [4] is defined as follows. In the encoding stage, the databases use generalized Pauli operators $\mathsf{X}(a)=\sum_{j=0}^{q-1}\ket{j+a}\bra{j}$ |, and $\mathsf{Z}(a)=\sum_{j=0}^{q-1}\omega^{tr(aj)}\ket{j}\bra{j}$ |, where $q=p^{r}$ with $p$ being a prime number, $a\in\mathbb{F}_{q}$ and $\omega=\exp(2\pi i/p)$.
@@
Let $\mathtt{G}$ be a $2N\times N$ SSO matrix, i.e., $\mathtt{G^{t}JG}=0$, where $\mathtt{J}=\begin{bmatrix}0&\mathtt{I}_{N}\\ -\mathtt{I}_{N}&0\end{bmatrix}$, and $\mathtt{H}$ be a $2N\times N$ matrix with full column rank. If $[\mathtt{G}~{}\mathtt{H}]$ is a full-rank matrix, then $\mathtt{G}^{\prime}=[0_{N}~{}\mathtt{I}_{N}][\mathtt{G}~{}\mathtt{H}]^{-1}$ is a feasible $N$-sum box transfer matrix.
@@ 5.1 Regime 1: $X+M\geq\frac{N}{2}$ and $E\leq 2X+2M-N$
To retrieve the $\theta$th message, the user sends the following queries to the $n$th server
@@
where $\bm{e}_{\theta}$ is a vector of length $K$ with $1$ in the $\theta$th index and zero otherwise, and $\bm{Z}_{ij}$ are uniform independent random vectors of length $K$ each, chosen by the user.
@@
where $Z_{j}^{\prime}(i)$ are generated uniformly at random and shared among servers. Note that the answers from the $N-U$ servers, $i_{1},\ldots,i_{N-U}$, for one instance only of the classical scheme, can be written compactly as
@@
$\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\mathcal{A}_{n}\left(\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\right)^{\dagger},\quad n\in[N]\setminus\mathcal{U}.$ (33)
@@
$\mathtt{X}(u_{n}\Delta^{\prime}_{n}(1))\mathtt{Z}(v_{n}\Delta^{\prime}_{n}(2))\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\mathcal{A}_{n}\left(\mathtt{X}(u_{n}\Delta^{\prime}_{n}(1))\mathtt{Z}(v_{n}\Delta^{\prime}_{n}(2))\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\right)^{\dagger}$
$=\mathtt{X}(u_{n}(A_{n}(1)+\Delta^{\prime}_{n}(1)))\mathtt{Z}(v_{n}(A_{n}(2)+\Delta^{\prime}_{n}(2)))\mathcal{A}_{n}\left(\mathtt{X}(u_{n}(A_{n}(1)+\Delta^{\prime}_{n}(1)))\mathtt{Z}(v_{n}(A_{n}(2)+\Delta^{\prime}_{n}(2)))\right)^{\dagger},$ (34)
@@
$\begin{bmatrix}\bm{I}_{1}(1)\\ \bm{I}_{1}(2)\\ \bm{W}_{\theta}(1)\\ \bm{W}_{\theta}(2)\\ \bm{I}_{2}(1)\\ \bm{I}_{2}(2)\\ \bm{\Delta}_{U}^{\prime}(1)\\ \bm{\Delta}_{U}^{\prime}(2)\end{bmatrix},$ (35)
where $\mathtt{\Lambda}_{1}=\mathtt{GRS}_{N,X+M}^{\bm{\alpha}_{[N]},\bm{u}_{[N]}}$, $\mathtt{\Lambda}_{2}=\mathtt{GRS}_{N,X+M}^{\bm{\alpha}_{[N]},v_{[N]}}$, $\mathtt{\Gamma}_{1}=\mathtt{GC}_{N,L}^{\bm{\alpha}_{[N]},\bm{u}_{[N]},\bm{f}_{[L]}}$, $\mathtt{\Gamma}_{2}=\mathtt{GC}_{N,L}^{\bm{\alpha}_{[N]},v_{[N]},\bm{f}_{[L]}}$, $a_{1}=\lceil\frac{N}{2}\rceil+1$, $a_{2}=\lfloor\frac{N}{2}\rfloor+1$, and $\mathtt{\Omega}=[\bm{e}_{n_{1}},\bm{e}_{n_{2}},\ldots,\bm{e}_{n_{U}}]$, where $\bm{e}_{n_{i}}$ is the unit vector of length $N$ and 00 in all locations except the $n_{i}$th index, and $\mathcal{U}=\{n_{1},\ldots,n_{U}\}$.
@@ 5.4 Regime 4: $X+M<\frac{N}{2}$ and $U+E>X+M$
In this setting, the usage of the classical scheme is preferable compared to the quantum scheme since there are not enough dimensions for transmitting message symbols using the quantum scheme or the rate of the classical scheme is better than the quantum scheme. For the sake of completeness, we present the classical scheme here. The sub-packetization length is $L=N-X-M-U$ and the storage is as in (27) and queries by the user are as in (28). In addition, the answers are computed as
@@ 6.1 Regime 1: $H+M+B\geq\frac{N}{2}$ and $E\leq 2H+2M+2B-N$
where $\bm{e}_{\theta}$ is a vector of length $K$ with $1$ in the $\theta$th index and zero otherwise, and $Z_{ij}$ are uniform independent random vectors of length $K$ each, chosen by the user.
Let =$\mathcal{B}=\{n_{i_{1}},\ldots,n_{i_{B}}\}$ denote the indices of the Byzantine servers, and as in the previous section $\mathcal{U}$ denotes the indices of the unresponsive servers. Thus, the answer from the $n$th server, $n\in[N]\setminus\left(\mathcal{B}\cup\mathcal{U}\right)$, is given by
@@
The answers from the $N-U$ servers, $i_{1},\ldots,i_{N-U}$, for one instance, can be written collectively as shown in [37], i.e.,
@@
and $\bm{\Delta}_{B}$ is a vector of length $N-U$ with only $B$ non-zero elements.
@@
$\begin{bmatrix}\tilde{\bm{W}}_{\theta}(i)\\ \tilde{\bm{I}}_{1}(i)\\ \tilde{\bm{I}}_{2}(i)\\ \tilde{\bm{\Delta}}_{B}(i)\end{bmatrix}=\begin{bmatrix}\bm{W}_{\theta}(i)\\ \hat{\bm{I}}_{[X+M+B]}(i)\\ \bm{0}_{2B}\end{bmatrix}+\mathtt{CSA}_{N-U,L,N-U}^{-1}\bm{\Delta}_{B}(i),$ (75)
@@ 6.4 Regime 4: $H+M+B<\frac{N}{2}$ and $2B+U+E>H+M+B$
In this setting, the usage of the classical scheme is preferable since there are not enough dimensions for transmitting message symbols using the classical scheme or the rate provided by the classical scheme is higher than the quantum scheme. For the sake of completeness, we present the classical scheme here. The sub-packetization length is $L=N-H-M-U-3B$ and the storage can be written as in (27) and queries by the user as in (28). In addition, the answers are computed as
@@
Note that in this section we assumed that the Byzantine servers use Pauli $\mathsf{X}(\cdot)$, and $\mathsf{Z}(\cdot)$ operations when transmitting the answers. This can be extended in the same exact way that is used in [47].
@@ 7 Decoding the Byzantine Errors
In the previous sections, the schemes are analyzed such that the user retrieves the $N-U$ symbols after applying the measurements. However, as discussed earlier, those symbols are contaminated with Byzantine errors. In this section, the decoding technique after measurements is developed in order to perform error correction for the Byzantine errors. We will use the following definitions for ease of notation:
@@
In addition, since $\begin{bmatrix}C_{1,\mathcal{J}(1)},\ldots,C_{N-U,\mathcal{J}(1)}\end{bmatrix}^{t},\ldots,\begin{bmatrix}C_{1,\mathcal{J}(B)},\ldots,C_{N-U,\mathcal{J}(B)}\end{bmatrix}^{t}$ are orthogonal to exactly $N-U-B$ rows of $\mathtt{CSA}_{N-U,L,N-U}$, we have
@@
Now, assume that the real Byzantine servers are $\mathcal{J}_{2}$ and the user guess is $\mathcal{J}_{1}$. Since we have $2B$ symbols, we can do the following: First, the user estimates $\hat{\bm{\Delta}}_{B}$ using the first $B$ symbols and using the last $B$ symbols, the estimation is either confirmed or denied. Since $\mathtt{\Psi}(\mathcal{J}_{1})$ is invertible, the user gets the following estimate $\bm{\overline{\Delta}}_{B}=\mathtt{\Psi}(\mathcal{J}_{1})^{-1}\mathtt{\Psi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}$. Afterwards, the user checks if $\mathtt{\Phi}(\mathcal{J}_{1})\mathtt{\Psi}(\mathcal{J}_{1})^{-1}\mathtt{\Psi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}=\mathtt{\Phi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}$. If this is satisfied, then to remove the Byzantine effect from the $N-U$ servers, the user multiplies $\bm{\overline{\Delta}}_{B}$ by $\mathtt{D}(N-U,\mathcal{J}_{1})$ and,

### v2

@@ 1.2 Related Work
As for the QPIR literature, in [7, 35] the eavesdropper variation was studied and the rate regions were found. In [36], the quantum unresponsive servers variation was analyzed and the rate regions were given. Finally, the quantum Byzantine servers version was analyzed and the rate regions were found in [37]. More applications for the new $N$-sum box abstraction can be found in [38, 39, 40].
@@ 2 Problem Formulation
[…]
@@
Upon receiving all the answers, the user must be able to decode the required message $W_{\theta}$ after applying the quantum measurement operations on the received quantum answers, i.e.,
$H(W_{\theta}|Y,Q_{[N]})=0,$ (9)
where $Y$ is the result of the measurements on the received quantum answers.
@@ 3 Main Results
$R_{Q}=\begin{cases}2\big(\frac{N-H-M-3B-U}{N}\big),&N-U>H+M+B\geq\frac{N}{2},\ E\leq 2H+2M+2B-N,\\ 2\big(\frac{N-H-M-3B-U-\frac{\delta}{2}}{N}\big),&N-U-\frac{\delta}{2}>H+M+B\geq\frac{N}{2},\ E>2H+2M+2B-N,\\ \frac{N-6B-2U-E}{N},&H+M+B<\frac{N}{2},\ 2B+U+E<H+M+B,\ N>6B+2U+E,\\ \frac{N-H-3B-U-M}{N},&H+M+B<\frac{N}{2},\ 2B+U+E>H+M+B,\ N>H+3B+U+M,\end{cases}$ (13)
@@
$R_{Q}=\begin{cases}2\big(\frac{N-H-M-3B-U}{N}\big),&N-U>H+M+B\geq\frac{N}{2},\ E\leq 2H+2M+2B-N,\\ 2\big(\frac{N-H-M-3B-U-\frac{\delta}{2}}{N}\big),&N-U-\frac{\delta}{2}>H+M+B\geq\frac{N}{2},\ E>2H+2M+2B-N,\\ \frac{N-6B-2U-E}{N},&H+M+B<\frac{N}{2},\ 2B+U+E<H+M+B,\ N>6B+2U+E,\\ \frac{N-H-3B-U-M}{N},&H+M+B<\frac{N}{2},\ 2B+U+E>H+M+B,\ N>H+3B+U+2M,\end{cases}$ (14)
@@
[…]
@@ 4.1 Quantum Information Theory
[…]
$|,\displaystyle\mathtt{\rho}_{A}=\sum_{j}p_{j}\ket{\psi_{j}}\bra{\psi_{j}}$, (15)
@@ 4.2 $N$-Sum Box Abstraction
The encoding and decoding structure using the $N$-sum box abstraction introduced recently in [4] is defined as follows. In the encoding stage, the databases use generalized Pauli operators $|\mathsf{X}(a)=\sum_{j=0}^{q-1}\ket{j+a}\bra{j}$, and $|\mathsf{Z}(a)=\sum_{j=0}^{q-1}\omega^{tr(aj)}\ket{j}\bra{j}$, where $q=p^{r}$ with $p$ being a prime number, $a\in\mathbb{F}_{q}$ and $\omega=\exp(2\pi i/p)$.
@@
Let $\mathtt{G}$ be a $2N\times N$ SSO matrix, i.e., $\mathtt{G^{t}JG}=0$, where $\mathtt{J}=\begin{bmatrix}0&\mathtt{I}_{N}\\ -\mathtt{I}_{N}&0\end{bmatrix}$, and $\mathtt{H}$ be a $2N\times N$ matrix with full column rank. If $\mathtt{G}~\mathtt{H}]$ is a full-rank matrix, then $\mathtt{G}^{\prime}=[0_{N}~\mathtt{I}_{N}][\mathtt{G}~\mathtt{H}]^{-1}$ is a feasible $N$-sum box transfer matrix.
@@ 5.1 Regime 1: $X+M\geq\frac{N}{2}$ and $E\leq 2X+2M-N$
[…]
@@
[…]
@@
where $Z_{j}^{\prime}(i)$ are generated uniformly at random and shared among servers. Note that the answers from the NUN-U servers, $i_{1},\ldots,i_{N-U}$, for one instance only of the classical scheme, can be written compactly as
@@
$\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\rho_{n}\left(\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\right)^{\dagger},\quad n\in[N]\setminus\mathcal{U},$ (33)
where $\rho_{n}$ is the density matrix for the quantum system $\mathcal{A}_{n}$.
@@
$\mathtt{X}(u_{n}\Delta^{\prime}_{n}(1))\mathtt{Z}(v_{n}\Delta^{\prime}_{n}(2))\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2)){\rho_{n}}\left(\mathtt{X}(u_{n}\Delta^{\prime}_{n}(1))\mathtt{Z}(v_{n}\Delta^{\prime}_{n}(2))\mathtt{X}(u_{n}A_{n}(1))\mathtt{Z}(v_{n}A_{n}(2))\right)^{\dagger}$
$=\mathtt{X}(u_{n}(A_{n}(1)+\Delta^{\prime}_{n}(1)))\mathtt{Z}(v_{n}(A_{n}(2)+\Delta^{\prime}_{n}(2))){\rho_{n}}\left(\mathtt{X}(u_{n}(A_{n}(1)+\Delta^{\prime}_{n}(1)))\mathtt{Z}(v_{n}(A_{n}(2)+\Delta^{\prime}_{n}(2)))\right)^{\dagger},$ (34)
@@
$\quad\begin{bmatrix}\bm{I}_{1}(1)\\ \bm{I}_{1}(2)\\ \bm{W}_{\theta}(1)\\ \bm{W}_{\theta}(2)\\ \bm{I}_{2}(1)\\ \bm{I}_{2}(2)\\ \bm{\Delta}_{U}^{\prime}(1)\\ \bm{\Delta}_{U}^{\prime}(2)\end{bmatrix}$, (35)
where $\mathtt{\Lambda}_{1}=\mathtt{GRS}_{N,X+M}^{\bm{\alpha}_{[N]},\bm{u}_{[N]}}$, $\mathtt{\Lambda}_{2}=\mathtt{GRS}_{N,X+M}^{\bm{\alpha}_{[N]},v_{[N]}}$, $\mathtt{\Gamma}_{1}=\mathtt{GC}_{N,L}^{\bm{\alpha}_{[N]},\bm{u}_{[N]},\bm{f}_{[L]}}$, $\mathtt{\Gamma}_{2}=\mathtt{GC}_{N,L}^{\bm{\alpha}_{[N]},v_{[N]},\bm{f}_{[L]}}$, $a_{1}=\lceil\frac{N}{2}\rceil+1$, $a_{2}=\lfloor\frac{N}{2}\rfloor+1$, and $\mathtt{\Omega}=[\bm{e}_{n_{1}},\bm{e}_{n_{2}},\ldots,\bm{e}_{n_{U}}]$, where $\bm{e}_{n_{i}}$ is the unit vector of length $N$ and 0 in all locations except the $n_{i}$th index, and $\mathcal{U}=\{n_{1},\ldots,n_{U}\}$.
@@ 5.4 Regime 4: $X+M<\frac{N}{2}$ and $U+E>X+M$
In this setting, the usage of the classical scheme is preferable compared to the quantum scheme since there are not enough dimensions for transmitting message symbols using the quantum scheme or the rate of the classical scheme is better than the quantum scheme. For the sake of completeness, we present the classical scheme here. The sub-packetization length is L=NXMUL=N-X-M-U and the storage is as in (27) and queries by the user are as in (28). In addition, the answers are computed as
@@ 6.1 Regime 1: $H+M+B\geq\frac{N}{2}$ and $E\leq 2H+2M+2B-N$
[…]
Let $\mathcal{B}=\{n_{i_{1}},\ldots,n_{i_{B}}\}$ denote the indices of the Byzantine servers, and as in the previous section $\mathcal{U}$ denotes the indices of the unresponsive servers. Thus, the answer from the $n$th server, $n\in[N]\setminus\left(\mathcal{B}\cup\mathcal{U}\right)$, is given by
@@
The answers from the NUN-U servers, $i_{1},\ldots,i_{N-U}$, for one instance, can be written collectively as shown in [37], i.e.,
@@
and $\bm{\Delta}_{B}$ is a vector of length NUN-U with only $B$ non-zero elements.
@@ 6.4 Regime 4: $H+M+B<\frac{N}{2}$ and $2B+U+E>H+M+B$
In this setting, the usage of the classical scheme is preferable since there are not enough dimensions for transmitting message symbols using the classical scheme or the rate provided by the classical scheme is higher than the quantum scheme. For the sake of completeness, we present the classical scheme here. The sub-packetization length is L=NHMU3BL=N-H-M-U-3B and the storage can be written as in (27) and queries by the user as in (28). In addition, the answers are computed as
@@
Note that in this section we assumed that the Byzantine servers use Pauli $\mathsf{X}(\cdot)$, and $\mathsf{Z}(\cdot)$ operations when transmitting the answers. This can be extended in the same exact way that is used in [46]. To explain intuitively, Kraus operators of any quantum operation that output Hilbert space is the same as the input Hilbert space can be decomposed into a linear combination of Weyl operators $\mathtt{W}(\bm{v})$, and any quantum code that corrects errors with Kraus operators $\{E_{i}\}_{i}$ can also correct errors with $\{F_{i}\}$ where $F_{i}=\sum_{j}m_{ij}E_{i}$. For each term individually, the scheme has the same rate thus by using superposition the rate is the same as considering the Weyl operators only being applied by the Byzantine servers.
@@ 7 Decoding the Byzantine Errors
In the previous sections, the schemes are analyzed such that the user retrieves the NUN-U symbols after applying the measurements. However, as discussed earlier, those symbols are contaminated with Byzantine errors. In this section, the decoding technique after measurements is developed in order to perform error correction for the Byzantine errors. We will use the following definitions for ease of notation:
@@
In addition, since $\begin{bmatrix}C_{1,\mathcal{J}(1)},\ldots,C_{N-U,\mathcal{J}(1)}\end{bmatrix}^{t},\ldots,\begin{bmatrix}C_{1,\mathcal{J}(B)},\ldots,C_{N-U,\mathcal{J}(B)}\end{bmatrix}^{t}$ are orthogonal to exactly NUBN-U-B rows of $\mathtt{CSA}_{N-U,L,N-U}$, we have
@@
Now, assume that the real Byzantine servers are $\mathcal{J}_{2}$ and the user guess is $\mathcal{J}_{1}$. Since we have $2B$ symbols, we can do the following: First, the user estimates $\hat{\bm{\Delta}}_{B}$ using the first $B$ symbols and using the last $B$ symbols, the estimation is either confirmed or denied. Since $\mathtt{\Psi}(\mathcal{J}_{1})$ is invertible, the user gets the following estimate $\bm{\overline{\Delta}}_{B}=\mathtt{\Psi}(\mathcal{J}_{1})^{-1}\mathtt{\Psi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}$. Afterwards, the user checks if $\mathtt{\Phi}(\mathcal{J}_{1})\mathtt{\Psi}(\mathcal{J}_{1})^{-1}\mathtt{\Psi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}=\mathtt{\Phi}(\mathcal{J}_{2})\hat{\bm{\Delta}}_{B}$. If this is satisfied, then to remove the Byzantine effect from the NUN-U servers, the user multiplies $\bm{\overline{\Delta}}_{B}$ by $\mathtt{D}(N-U,\mathcal{J}_{1})$ and,